# AGRAG: Advanced Graph-based Retrieval-Augmented Generation for LLMs

### v1

@@ Abstract
Graph-based retrieval-augmented generation (Graph-based RAG) has demonstrated significant potential in enhancing Large Language Models (LLMs) with structured knowledge. However, existing methods face three critical challenges: Inaccurate Graph Construction, caused by LLM hallucination; Poor Reasoning Ability, caused by failing to generate explicit reasons telling LLM why certain chunks were selected; and Inadequate Answering, which only partially answers the query due to the inadequate LLM reasoning, making their performance lag behind NaiveRAG on certain tasks. To address these issues, we propose AGRAG, an advanced graph-based retrieval-augmented generation framework. When constructing the graph, AGRAG substitutes the widely used LLM entity extraction method with a statistics-based method, avoiding hallucination and error propagation. When retrieval, AGRAG formulates the graph reasoning procedure as the Minimum Cost Maximum Influence (MCMI) subgraph generation problem, where we try to include more nodes with high influence score, but with less involving edge cost, to make the generated reasoning paths more comprehensive. We prove this problem to be NP-hard, and propose a greedy algorithm to solve it. The MCMI subgraph generated can serve as explicit reasoning paths to tell LLM why certain chunks were retrieved, thereby making the LLM better focus on the query-related part contents of the chunks, reducing the impact of noise, and improving AGRAG’s reasoning ability. Furthermore, compared with the simple tree-structured reasoning paths, our MCMI subgraph can allow more complex graph structures, such as cycles, and improve the comprehensiveness of the generated reasoning paths.
@@ I Introduction
Figure 1: An overview of Graph-based RAG models. It first extracts entities and relations as graph nodes and edges. When retrieval, they will map query to graph nodes, and utilizing the graph structure to support query answering.
Large Language Models (LLMs) [1, 2, 3, 4, 5] have demonstrated remarkable capabilities in natural language understanding and generation [6, 7, 8, 9, 10, 11, 12, 13, 14]. However, their reliance on static, internalized knowledge often leads to factual hallucinations and limited adaptability to dynamic information [15, 16, 17, 18], and retraining or finetuning the LLM can lead to high computational costs [19, 20]. Retrieval-Augmented Generation (RAG) addresses this by integrating external factual knowledge during inference, improving accuracy and enabling real-time updates. Yet, traditional RAG methods’ retrieval [21, 15] is based on simplistic mechanisms, such as the whole passages’ dense vector similarity, make them overlook the capture of more fine-grained semantic relationships between entities [22, 23, 24]. To overcome these limitations, researchers have proposed various Graph-based RAG models [25, 26, 27, 28, 29, 30, 31]. The framework of Graph-based RAG models, is shown in Figure 1. These models first utilize LLMs to extract entities and relations from the source documents to serve as graph nodes and edges and construct a graph, and then leverage this graph-structured knowledge to model complex relationships between entities to enable structured reasoning over inter-connected information, enabling the capture of more fine-grained entity-level relationships when retrieval. [22, 24].
Currently, the Graph-based RAG models can be classified into two categories, i.e., community-based models, and walking-based models. The community-based models [25, 28, 26] first map the query to graph nodes considering their semantic similarity or keyword matching. Then, they will retrieve the mapped graph nodes’ one or multi-hop neighborhood. On the other hand, with similar mapping approach from query to graph nodes, walking-based models [29, 32, 27, 30, 31] apply walk-based graph algorithms, such as Personalized PageRank (PPR) [33], based on the mapped graph nodes, to score graph nodes and edges based on the graph’s topological structure and query information, and then retrieve Top-k text chunks based on high-scoring graph nodes, capturing the fine-grained entity level relationships.
@@
    Inaccurate Graph Construction: Current Graph-based RAG models apply LLM for entity and relation extraction to construct graphs. However, the LLM can inevitably suffer from hallucination issue [15, 16, 17, 18] and leading to inaccurate entity extraction, introducing noise into the constructed graph and propagate through the whole pipeline, affecting the model effectiveness.
@@
    Inadequate Answer: Although we can simply combine the existing RAG models with the current shortest path-based graph algorithm to provide explicit reasoning chains to aid LLM query answering, improve the model’s reasoning ability. However, on tasks that require summarization of current knowledge [34, 35, 36], this simple approach can lead to suboptimal performance, as the shortest path on the graph can be overly simple, and can omit many query-related facts, leading to incomplete reasoning chains, which makes the model tend to generate inadequate query answer with low information coverage and faithfulness (See Table VI and subsection IV-D).
To address these issues, we propose a novel Advanced Graph-based Retrieval-Augmented Generation framework, called AGRAG. We divide the AGRAG framework into three steps: In Step 1 Data Preparation, AGRAG first extracts entities with a TFIDF [37] based entity extraction method, avoiding LLM hallucination, and resulting in better graph construction accuracy. Then, AGRAG applies LLM to detect the possible relation w.r.t. the extracted entity and the text chunk, formulating entity-relation-entity triples, namely knowledge triple. Then these knowledge triples, as well as the text chunks involved, together formulate a Knowledge Graph (KG). In Step 2 Graph Retrieval, AGRAG maps the query to knowledge triples in KG based on their semantic similariy, then calculate the PPR score [33] as node influence based on the KG structure and the mapped triples. In the meantime, AGRAG also calculates edge weights based on the semantic similarity between the query and the edge’s corresponding triple. After that, AGRAG generates a Minimum Cost and Maximum Influence (MCMI) reasoning subgraph via. a greedy algorithm, which is calculated based on the shortest spanning tree of the mapped triples. It tries to maximize the node influence score, while minimize the edge cost in the expansion of each node, in order to avoid introducing noise. Compared to other shortest path-based algorithms, the MCMI subgraph allows for more complex structures, such as rings, which results in reasoning chains containing more query-related entities and relations, making them more comprehensive. This, in turn, helps the model perform better on complex tasks that require summarization, resulting in better reasoning ability and more comprehensive query answers. In Step 3 Hybrid Retrieval, AGRAG carries out hybrid retrieval based on the average of dense representation vector similarity score and sparse BM25 [38] keyword matching score between query and text chunks, providing extra contexts for query answering. At last, AGRAG will input the resulting subgraph from the Graph Retrieval step and chunks from hybrid retrieval to LLM for final query answering.
@@
  • We present the AGRAG framework, which consists of three steps: Data Preparation, Graph Retrieval, and Hybrid Retrieval, aiming at addressing the Inaccurate Graph Construction, Poor Reasoning Ability, and Inadequate Answer issues of current Graph-based RAG models.
@@
  • According to our experiments, AGRAG achieve better effectiveness and efficiency compared with current state-of-the-art RAG models on six tested tasks.
@@ II-A Traditional RAG Models
Although LLMs are inherently capable of inference without fine-tuning [1, 2, 3, 4, 5]. However, the lack of fine-tuning with up-to-dated and domain specific knowledge can lead LLMs to generate incorrect answers, as they lack task-specific knowledge, these incorrect answers are often referred to as hallucinations [59]. To mitigate hallucinations, researchers try to directly retrieve text documents outside of LLM based on the query as side information to help the LLM query answering [60, 61, 62, 63, 64, 65, 66, 67], lead to the emergence of early RAG models, namely traditional RAG models. The earliest LLM-RAG model is the simple NaiveRAG [15], when indexing, it split the texts into chunks, and encode them to semantic vector representations. When retrieval, the model retrieve several text chunks with top semantic similarity to the query.
Based on NaiveRAG, researchers have walking-based various models to further tailor the text chunks and improve the retrieval quality. For example, Propositionizer [60] applies a fine-tuned LLM to convert the text chunks into atomic expressions, namely propositions, to facilitate fine-grained information retrieval. Prompt compressor models [68, 69], such as LongLLMLingua [68] apply LLM’s generation perplexity on to filter out un-important tokens in the text chunks, try to maintain a more concise retrieval result for LLM input. However, the retrieval procedures of these models remain unstructured and struggle to capture the nuanced semantic and structured relationship within the text chunks, which impedes the effectiveness of LLMs [15, 16, 17, 18].
@@ II-B Graph-based RAG Models
To overcome the aforementioned limitations of traditional RAG models, researchers have begun to develop Graph-based RAG models [25, 26, 27, 28, 29, 30, 31]. These models usually extract graph components with LLMs from text corpis to index them into graphs. When retrieval, these models will utilize the graph topology structure to capture complex relationships between entities within and across text chunks.
II-B1 Community-based Models
Early Graph-based RAG models typically retrieve specific neighborhoods from the graph based on the input query. The pioneering work in this category is GraphRAG [25], as illustrated in Figure 1, which constructs a graph and aggregates nodes into communities, followed by generating community summaries to assist the LLM in making predictions. However, GraphRAG requires re-computing the graph communities after each index update, leading to significant computational overhead.
To address this limitation, LightRAG [26] eliminates the explicit detection of graph communities, as well as the computation of community summaries. Instead, it retrieves the most semantically similar entities along with their 1-hop graph neighborhoods and the corresponding text passages containing these entities as external knowledge, which are then fed to the LLM. Nevertheless, this approach can still incur computational costs and may lead to hallucinations, as the 1-hop neighborhoods can be large and often contain noisy information.
Moreover, all these community-based models are limited in their ability to capture structural information beyond a small local neighborhood, thereby restricting their ability for long-range reasoning.
@@ II-B1 Community-based Models
$\mathcal{T}$, $\mathcal{T}_{c}$ Original texts and splited text chunks.
@@
$d$, $\mathbf{p}$, $\mathbf{P}$ The damping factor, personalization vector and transition probablity matrix of PPR.
$\mathcal{N}(\cdot)$ 1-hop neighbor set of a node in KG.
$r_{MCMI}$ MCMI subgraph’s average score-cost ratio.
@@
II-B2 Walking-based Models
To further enhance the long-range graph reasoning capabilities of community-based models, researchers have proposed walking-based models [29, 32, 27, 31], which apply walking-based graph algorithms, to detect and capture global topological information within the Knowledge Graphs (KGs).
Among these, HippoRAG [29] and Fast-GraphRAG [32] utilize the Personalized PageRank (PPR) algorithm to traverse the constructed KG and assign relevance scores to graph nodes with respect to the input query. During retrieval, these models select text chunks based on the aggregated PPR scores of the graph nodes they contain. StructRAG [27] trains a router module that learns to identify the optimal reasoning path over the graph based on the input query. HippoRAG2 [31] further improves upon HippoRAG by incorporating text chunks as passage nodes directly into the KG. It retrieves text chunks based on the individual PPR scores of their corresponding passage nodes within the graph.
@@ III Methodology
In this section, we first introduce the overall AGRAG framework, then detail its three core steps: Data Preparation, Graph Retrieval, and Hybrid Retrieval. Finally, we discuss the computational hardness and complexity analysis of AGRAG.
Figure 2: An overview of AGRAG. In Step 1, AGRAG constructs a KG based on the text corpus. In Step 2, the query is mapped to relevant triplet facts in KG, then node score and edge weights are assigned based on PPR algorithm and the semantic similarity between query and each fact, respectively. In Step 3, AGRAG generates a MCMI subgraph, providing explicit reasoning chains in addition the most semantically similar text chunks by Hybrid Retrieval (HR) to support LLM query answering.
@@ III-A Framework Overview
Step 1: Data Preparation. In this step, AGRAG constructs a Knowledge Graph (KG) $\mathcal{KG}(\mathcal{V},\mathcal{E})$ based on the given text corpus $\mathcal{T}$. Here, we substitute other Graph-based RAG baselines’ LLM entity extraction to a statistics-based method based on TFIDF [37], and extract entities from text chunks in $\mathcal{T}_{c}$ that splitted from the texts in corpus $\mathcal{T}$. This method can avoid LLM hallucination, and hence reduce the noise in the constructed graph, leading to better effectiveness. After entity extraction, we apply LLM to detect relation edges w.r.t. the extracted entity nodes and the text chunk. Furthermore, we also construct text chunks into KG as passage nodes, and link them to all their extracted entity nodes with node-to-passage edges. The detailed pseudo code of Step 1 is shown in algorithm 1, and more details can be found in Section III-B.
Step 2: Graph Retrieval. In this step, we firstly map the query $q$ to KG fact triples considering their semantic similarities, after that, we calculate the PPR score [33] based on these mapped fact triples $F$ as the node influence score $\mathcal{S}_{\mathcal{V}}$. On the other hand, we calculate the edge costs $\mathcal{C}_{\mathcal{E}}$ considering the semantic similarity between query and facts’ vector representations. the higher their similarity is, the lower the cost is. Compared with solely applying PPR to score nodes for retrieval, the edge cost further provides constraints on the involved nodes based on the semantic correspondence between query and facts, removing the need for the human-defined PPR retrieval threshold. After that, we initialize the MCMI subgraph with the generated MCST, and apply our greedy algorithm by iteratively expanding the MCMI subgraph by selecting its neighboring node with the highest influence score-edge cost ratio (i.e., $s/c$). If the highest ratio is smaller than the average ratio of the current MCMI, the iteration stops and the current MCMI subgraph is returned. We prove the MCMI generation problem to be NP-hard, and our greedy algorithm in algorithm 3 can achieve 2-approximate. Please refer to Section III-C for further details.

### v2

@@ Abstract
Graph-based retrieval-augmented generation (Graph-based RAG) has demonstrated significant potential in enhancing Large Language Models (LLMs) with structured knowledge. However, existing methods face three critical challenges: Inaccurate Graph Construction, caused by LLM hallucination; Poor Reasoning Ability, caused by failing to generate explicit reasons telling LLM why certain chunks were selected; and Inadequate Answering, which only partially answers the query due to the inadequate LLM reasoning, making their performance lag behind NaiveRAG on certain tasks. To address these issues, we propose AGRAG, an advanced graph-based retrieval-augmented generation framework. When constructing the graph, AGRAG substitutes the widely used LLM entity extraction method with a statistics-based method, avoiding hallucination and error propagation. During retrieval, AGRAG formulates the graph reasoning procedure as the Minimum Cost Maximum Influence (MCMI) subgraph generation problem, where we try to include more nodes with high influence score, but with less involving edge cost, to make the generated reasoning paths more comprehensive. We prove this problem to be NP-hard, and propose a greedy algorithm to solve it. The MCMI subgraph generated can serve as explicit reasoning paths to tell LLM why certain chunks were retrieved, thereby making the LLM better focus on the query-related part contents of the chunks, reducing the impact of noise, and improving AGRAG’s reasoning ability. Furthermore, compared with the simple tree-structured reasoning paths, our MCMI subgraph can allow more complex graph structures, such as cycles, and improve the comprehensiveness of the generated reasoning paths. The code and prompt of AGRAG are released at: https://github.com/Wyb0627/AGRAG.
@@ I Introduction
Large Language Models (LLMs) [63, 33, 65, 1, 79] have demonstrated remarkable capabilities in natural language understanding and generation [80, 10, 47, 36, 5, 87, 6, 49, 53]. However, their reliance on static, internalized knowledge often leads to factual hallucinations and limited adaptability to dynamic information [22, 56, 72, 46], and retraining or finetuning the LLM can lead to high computational costs [51, 19]. Retrieval-Augmented Generation (RAG) addresses this by integrating external factual knowledge during inference, improving accuracy and enabling real-time updates. Yet, traditional RAG methods’ retrieval [40, 22] is based on simplistic mechanisms, such as whole passage’s dense vector similarity, making them overlook the capture of more fine-grained semantic relationships between entities [26]. To overcome these limitations, researchers have proposed various Graph-based RAG models [17, 24, 42, 29, 35, 43, gutiérrez2025ragmemory, 69, 89]. These models first utilize LLMs to extract entities and relations from the source documents to serve as graph nodes and edges and construct a graph, and then leverage this graph-structured knowledge to model complex relationships between entities to enable structured reasoning over inter-connected information, enabling the capture of more fine-grained entity-level relationships. [26].
Currently, the Graph-based RAG models can be classified into two categories, i.e., community-based models, and walking-based models. The community-based models [17, 29, 24] first map the query to graph nodes considering their semantic similarity or keyword matching. Then, they will retrieve the mapped graph nodes’ one or multi-hop neighborhood. On the other hand, with similar mapping approach from query to graph nodes, walking-based models [35, 13, 42, 43, gutiérrez2025ragmemory] apply walk-based graph algorithms, such as Personalized PageRank (PPR) [27], based on the mapped graph nodes, to score graph nodes and edges based on the graph’s topological structure and query information, and then retrieve Top-k text chunks based on high-scoring graph nodes, capturing the fine-grained entity level relationships.
@@
    Inaccurate Graph Construction: Current Graph-based RAG models apply LLM for entity and relation extraction to construct graphs. However, the LLM can inevitably suffer from the hallucination issue [22, 56, 72, 46], leading to inaccurate entity extraction, introducing noise into the constructed graph and propagate through the whole pipeline, affecting the model effectiveness.
@@
    Inadequate Answer: Although we can simply combine the existing RAG models with the current shortest path-based graph algorithm to provide explicit reasoning chains to aid LLM query answering, improve the model’s reasoning ability. However, on tasks that require summarization of current knowledge [74, 25, 89], this simple approach can lead to suboptimal performance, as the shortest path on the graph can be overly simple, and can omit many query-related facts, leading to incomplete reasoning chains, which makes the model tend to generate inadequate query answer with low information coverage and faithfulness (See Table VI and Section IV-D).
To address these issues, we propose a novel Advanced Graph-based Retrieval-Augmented Generation framework, called AGRAG. We divide the AGRAG framework into three steps: In Step 1 Data Preparation, AGRAG first extracts entities with a TFIDF [64] based entity extraction method, avoiding LLM hallucination, and resulting in better graph construction accuracy. Then, AGRAG applies LLM to detect the possible relation w.r.t. the extracted entity and the text chunk, formulating entity-relation-entity triples, namely knowledge triples. Then these knowledge triples, as well as the text chunks involved, together formulate a Knowledge Graph (KG). In Step 2 Graph Retrieval, AGRAG maps the query to knowledge triples in KG based on their semantic similarity, then calculate the PPR score [27] as node influence based on the KG structure and the mapped triples. In the meantime, AGRAG also calculates edge weights based on the semantic similarity between the query and the edge’s corresponding triple. After that, AGRAG generates a Minimum Cost and Maximum Influence (MCMI) reasoning subgraph via a greedy algorithm, which is calculated based on the shortest spanning tree of the mapped triples. It tries to maximize the node influence score, while minimize the edge cost in the expansion of each node, in order to avoid introducing noise. Compared to other shortest path-based algorithms, the MCMI subgraph allows for more complex structures, such as rings, which results in reasoning chains containing more query-related entities and relations, making them more comprehensive. This, in turn, helps the model perform better on complex tasks that require summarization, resulting in better reasoning ability and more comprehensive query answers. In Step 3 Hybrid Retrieval, AGRAG carries out hybrid retrieval based on the average of dense representation vector similarity score and sparse BM25 [59] keyword matching score between query and text chunks, providing extra contexts for query answering. At last, AGRAG will input the resulting subgraph from the Graph Retrieval step and chunks from hybrid retrieval to LLM for final query answering.
@@
  • We present the AGRAG framework, which consists of three steps: Data Preparation, Graph Retrieval, and Hybrid Retrieval, aiming to address the Inaccurate Graph Construction, Poor Reasoning Ability, and Inadequate Answer issues of current Graph-based RAG models.
@@
  • According to our experiments, AGRAG achieves better effectiveness and efficiency compared with current state-of-the-art RAG models on six tested tasks.
@@ II-A Traditional RAG Models
Although LLMs are inherently capable of inference without fine-tuning [63, 33, 65, 1, 79], the lack of fine-tuning with up-to-date and domain specific knowledge can lead LLMs to generate incorrect answers. As they lack task-specific knowledge, these incorrect answers are often referred to as hallucinations [83]. To mitigate hallucinations, researchers try to directly retrieve text documents outside of LLM based on the query as side information to help the LLM query answering [12, 60, 30, 71, 8, 54, 76, 9, 88, 70], lead to the emergence of early RAG models, namely traditional RAG models. The earliest LLM-RAG model is the simple NaiveRAG [22], when indexing, it splits the texts into chunks, and encodes them to semantic vector representations. When retrieval, the model retrieves several text chunks with top semantic similarity to the query.
Based on NaiveRAG, researchers have proposed various walking-based models to further tailor the text chunks and improve the retrieval quality. For example, Propositionizer [12] applies a fine-tuned LLM to convert the text chunks into atomic expressions, namely propositions, to facilitate fine-grained information retrieval. Prompt compressor models [55, 34], such as LongLLMLingua [55] apply LLM’s generation perplexity to filter out unimportant tokens in the text chunks, try to maintain a more concise retrieval result for LLM input. However, the retrieval procedures of these models remain unstructured and struggle to capture the nuanced semantic and structured relationship within the text chunks, which impedes the effectiveness of LLMs [22, 56, 72, 46].
@@ II-B Graph-based RAG Models
To overcome the aforementioned limitations of traditional RAG models, researchers have begun to develop Graph-based RAG models [17, 24, 42, 29, 35, 43, gutiérrez2025ragmemory]. These models usually extract graph components with LLMs from text corpora to index them into graphs. When retrieval, these models will utilize the graph topology structure to capture complex relationships between entities within and across text chunks.
II-B1 Tree-based Models
To address the limitation of flat retrieval in capturing long-range dependencies and thematic information, tree-based models organize text chunks into hierarchical structures. RAPTOR [60] recursively clusters and summarizes text chunks to construct a tree from the bottom up, enabling the retrieval of information at varying levels of abstraction to answer high-level thematic queries. Building on this, SiReRAG [82] constructs dual index trees to capture both semantic similarity and entity-based relatedness. It builds a similarity tree via recursive summarization and a relatedness tree that groups fine-grained propositions sharing the same entities, thereby indexing both similar and related information to support complex multihop reasoning.
However, these tree-based models can only generate tree-structured reasoning chains, which restricts their adaptivity to different tasks.
II-B2 Community-based Models
Early Graph-based RAG models typically retrieve specific neighborhoods from the graph based on the input query. The pioneering work in this category is GraphRAG [17], it aggregates nodes into communities and generates community summaries to encapsulate global information. To reduce the computational overhead of community detection, LightRAG [24] retrieves semantically similar entities along with their 1-hop graph neighborhoods and corresponding text passages, eliminating the need for pre-computed summaries.
Moving beyond fixed neighborhoods, G-Retriever [28] formulates subgraph retrieval as a Prize-Collecting Steiner Tree (PCST) optimization problem, identifying connected subgraphs that maximize relevance while minimizing costs. Recently, MIXRAG [45] introduces a Mixture-of-Experts (MoE) framework to handle diverse query intents, dynamically routing queries to specialized entity, relation, or subgraph experts and employing a query-aware graph encoder to filter noise.
Moreover, all these community-based models can be limited in capturing information beyond a small local neighborhood, thereby restricting their long-hop reasoning ability [48, 86].
@@ II-B2 Community-based Models
$\mathcal{T}$, $\mathcal{T}_{c}$ Original texts and split text chunks.
@@
$d$, $\mathbf{p}$, $\mathbf{P}$ The damping factor, personalization vector and transition probability matrix of PPR.
$\mathcal{N}(\cdot)$ 1-hop neighbor set of node or subgraph in KG.
$r_{MCMI}$ MCMI subgraph’s total cost-score ratio.
@@
II-B3 Walking-based Models
To further enhance the long-range graph reasoning capabilities of community-based models, researchers have proposed walking-based models [35, 13, 42, gutiérrez2025ragmemory], which apply walking-based graph algorithms, to detect and capture global topological information within the Knowledge Graphs (KGs).
Among these, HippoRAG [35] and Fast-GraphRAG [13] utilize the Personalized PageRank (PPR) algorithm to traverse the constructed KG and assign relevance scores to graph nodes with respect to the input query. During retrieval, these models select text chunks based on the aggregated PPR scores of the graph nodes they contain. StructRAG [42] trains a router module that learns to identify the optimal reasoning path over the graph based on the input query. HippoRAG2 [gutiérrez2025ragmemory] further improves upon HippoRAG by incorporating text chunks as passage nodes directly into the KG. It retrieves text chunks based on the individual PPR scores of their corresponding passage nodes within the graph.
@@ III Methodology
In this section, we first introduce the overall AGRAG framework, then detail its three core steps: Data Preparation, Graph Retrieval, and Hybrid Retrieval.
Figure 1: An overview of AGRAG. In Stage 1, AGRAG constructs a KG based on the text corpus. In Stage 2, the query is mapped to relevant triplet facts in KG, then node score and edge weights are assigned based on PPR algorithm and the semantic similarity between query and each fact, respectively. In Stage 3, AGRAG generates an MCMI subgraph, providing explicit reasoning chains in addition to the most semantically similar text chunks by Hybrid Retrieval (HR) to support LLM query answering.
@@ III-A Framework Overview
Step 1: Data Preparation. In this step, AGRAG constructs a Knowledge Graph (KG) $\mathcal{KG}(\mathcal{V},\mathcal{E})$ based on the given text corpus $\mathcal{T}$. Here, we substitute other Graph-based RAG baselines’ LLM entity extraction to a statistics-based method based on TFIDF [64], and extract entities from text chunks in $\mathcal{T}_{c}$ that split from the texts in corpus $\mathcal{T}$. This method can avoid LLM hallucination, and hence reduce the noise in the constructed graph, leading to better effectiveness. After entity extraction, we apply LLM to detect relation edges w.r.t. the extracted entity nodes and the text chunk. Furthermore, we also construct text chunks into KG as passage nodes, and link them to all their extracted entity nodes with node-to-passage edges. The detailed pseudo code of Step 1 is shown in Algorithm 1, and more details can be found in Section III-B.
Step 2: Graph Retrieval. In this step, we firstly map the query $q$ to KG fact triples considering their semantic similarities, after that, we calculate the PPR score [27] based on these mapped fact triples $F$ as the node influence score $\mathbf{s}$. On the other hand, we calculate the edge costs $\mathcal{C}_{\mathcal{E}}$ considering the semantic similarity between query and facts’ vector representations. the higher their similarity is, the lower the cost is. Compared with solely applying PPR to score nodes for retrieval, the edge cost further provides constraints on the involved nodes based on the semantic correspondence between query and facts, removing the need for the human-defined PPR retrieval threshold. After that, we initialize the MCMI subgraph with the generated MCST, and apply our greedy algorithm by iteratively expanding the MCMI subgraph by selecting its neighboring node with the lowest edge cost-influence score ratio (i.e., $c/s$). If the lowest ratio is larger than the total ratio of the current MCMI, the iteration stops and the current MCMI subgraph is returned. We prove the MCMI generation problem to be NP-hard, and our greedy algorithm in Algorithm 3 can achieve a 2-approximation. Please refer to Section III-C for further details.